OpenStudy (anonymous):

find dy/dx of the (8th root of t)^t (t^(1/8))^t

OpenStudy (anonymous):

i keep getting ((t*(1/8)t^(-7/8))/(t^1/8))* y

OpenStudy (anonymous):

does anyone know how to do this?

OpenStudy (anonymous):

\[ \frac{d}{dt} t^{t/8} = {t}^{1/8\,t} \left( 1/8\,\ln \left( t \right) +1/8 \right) \]